RC meeting and Run Plan, Monday Nov 19
RC Meeting Information
RC Daily Meetings are at 4:30 PM, 2nd floor Counting House Meeting Room
- Bluejeans remote connection information:
To join the Meeting: https://bluejeans.com/861439920 To join via Room System: Video Conferencing System: bjn.vc -or-126.96.36.199 Meeting ID : 861439920 To join via phone : 1) Dial: +1.408.740.7256 (US (San Jose)) +1.888.240.2560 (US Toll Free)) +1.408.317.9253 (US (Primary, San Jose)) (see all numbers - http://bluejeans.com/numbers) 2) Enter Conference ID : 861439920
- Detailed Kinematic Group Settings
- KaonLT-SIDIS Group 2018-2019 Shift Schedule
- Hall-C Controls
- Hall-C Logbook
- MCC White Board
- Taking kinematic Group 13 (Q2 = 4.76, x = 0.55, pi-)
- RC Daily Update Saturday, November 17
- RC Daily Update Sunday, November 18
- RC Daily Update Monday, November 19
Current Run Plan
- Order of Kinematic Groups
Notices in Effect for Shift Crew
- Shift leaders, please fill out the check list once per shift!
- It is no longer required to fill out the standard.kinematics file online, it is now done automagically
- Split the large desired charge between multiple runs (~1 hour per run)
- Listed currents are estimates. For runs at less than 60uA, increase current until SHMS singles rates are between 400k-500k.
- How-to power cycle PLC I/O chassis: https://logbooks.jlab.org/entry/3625862, https://logbooks.jlab.org/entry/3625900
- Worsening of coincidence time resolution
- Cause of coin. time jitter (resolution loss) may have been found? -Brad S.
- Coincidence time after insertion of passive delay -Dave G.
- Follow-up Re: Coincidence time after insertion of passive delay -Dave G. with comment from Dave M.
- More jitter tests -- nothing found on scope... -Brad S.
- HMS/SHMS TOF resolutions using hodoscope beta -Mark J.
- RF time and coincidence time -Mark J.
- Yet more jitter tests -- found small issue w/ ref time, but nothing that impacts coin. times -Brad S.
Planned Accesses to the Hall
- Monday, Nov 26: aerogel change and HGC work needs to be coordinated with Walter K., we must be in restricted access for this work
Planned Accelerator Configuration Changes
- Tuesday, Nov 20: CHL UPS battery change, should be noninvasive however, induce CHL anomaly is possible
- Wednesday, Nov 21: Hall A goes offline and MCC wants to conduct high current beam study of energy spread at 40 uA. It is unclear that we will be effected
- Monday, Nov 26: Energy change to 1.86 GeV/pass, we will be moved over to 500 MHz laser.
- Harp scan at 3H07A and 3H07B with raster off to address possible issues with harp (https://logbooks.jlab.org/entry/3624463)
- Vijay Kumar (Regina group) has requested a walkthrough during the next access
- Jay B. would like to conduct another beam steering test (5 minute procedure) during the next polarity change. He requests to only be contacted if the polarity change occurs during the hours of 0630 and 1930.